Can anyone help me on a word problem, putting the situation into a formula, and solving? I am struggling on this one:
Great white sharks have triangular teeth with a height that is 1 centimeter longer than the base. If the area of one tooth is 15 square centimeters, find it's base and height.
I have played with formulas and keep scratching my head. Any help would be appreciated!

Answer by stanbon(75887)

Formula: Area=(1/2)(base)(height)
15=(1/2)b(b+1)
30=b^2+b
b^2+b-30=0
(b+6)(b-5)=0
base=5 cm so height=6 cm
Cheers,
Stan H.